solve the equation for x
25x^2=13

Answer by malakumar_kos@yahoo.com(315)


solve the equation for x
25x^2=13

25x^2 = 13

x^2 = 13/25
x = +-sq.rt 13/25
x = 1/5 (+-sq.rt 13)